If the price of a slice of pizza rises from $2.50 to $3, and quantity demanded falls from 10,000 slices to 7,400 slices, using the formula for arc price elasticity

i.    What is the percentage change in price?


1
Expert's answer
2021-07-22T10:00:36-0400
"E=\frac{Q_2-Q_1}{Q_2+Q_1}\times \frac{p_2+p_1}{p_2-p_1}"

"E=-1.61"